OpenStudy (anonymous):

2x^2+7x+6 2x^2+7x+6 @Mathematics

OpenStudy (anonymous):

i tried 6, 1

OpenStudy (anonymous):

= (2x + 3 )(x + 2 )

OpenStudy (anonymous):

thnx jimmyrep

OpenStudy (anonymous):

\[\huge x_1=\frac{-7+\sqrt{40-24}}{4}\] \[\huge x_1=\frac{-7-\sqrt{40-24}}{4}\]

OpenStudy (anonymous):

you need 3 numbers when multiplied give + 6 and when added give + 7 they are 2 and 3 and anotehr 2 to give 4 (2x2)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

2x^2+7x+6 2x^2 + 4x + 3x + 6 2x(x + 2) + 3(x + 2) (2x + 3)(x + 2)